OpenStudy (anonymous):

You just purchased a cellular phone and are trying to decide the best cellular phone company in which to give your business. When you contacted the Talks-A-Lot Company, they were offering a monthly plan of $40 for 500 minutes and $0.25 for each minute exceeding the 500 minutes. In the Sunday paper you see an ad for the Chat-Away Company, which offers a monthly plan of $35 for 500 minutes and $0.30 for each minute exceeding the 500 minutes. How many minutes would you have to talk over and above the 500 minutes for the cost to be the same with both companies? What would be the equal cost? 1

OpenStudy (anonymous):

100 minutes over 500 minutes

OpenStudy (cruffo):

Let T(x) represent the piecewise function describing the Talks-A-Lot plan. If x > 500, then \[T(x) = 40 + 0.25(x-500)\]

OpenStudy (anonymous):

equal cost will be $65

OpenStudy (anonymous):

do u need the working??

OpenStudy (anonymous):

yes i do

OpenStudy (cruffo):

Let C(x) represent the piecewise function describing the Chat-Away plan. If x > 500, then \[C(x) = 35 + 0.30(x-500).\] You want to find the value of x that makes T(x) = C(x). That part has been answered by Harkirat :)

OpenStudy (anonymous):

see both companies are giving 500 minutes for a fixed cost so extra cost will be for the extra minutes talked Let this time = xmin so for Talk-a-lot the cost will be = 40 + 0.25x for chat-away the cost will be = 35 + .30x since the two costs have to be same, 40+0.25x = 35+0.30x 40-35 = 0.30x - 0.25x 5=0.05x 5/0.05 = x 100 = x So the extra minutes used is 100 equal cost will be talk-a-lot = 40 + 0.25(100) = 40 + 25 = 65 chat away = 35 + 0.30(100) = 35 + 30 = 65

OpenStudy (anonymous):

thanks alot you guys this is really helping me alot :)